if one leg of a right triangle is 8 inches long, and the other leg is 12 inches long, how many inches long is the triangle’s hypotenuse?
Answer by unlockmath(1688)

Hello,
We know the formula is A^2 + B^2 = C^2
So plug in the values:
8^2 + 12^2 = C^2
64+ 144 = C^2
208=C^2
Square root both sides and we get:
c= +-sq rt 208
Make Sense?
RJ
www.math-unlock.com